A researcher wishes to know with 95% confidence the proportion of people who own a desktop computer. A previous study showed that 40% of those interviewed owned a desktop. The researcher wishes to be accurate within 2% of the true proportion.

Find the minimum sample size necessary.
Answer #1

SOLUTION:

From given data,

A researcher wishes to know with 95% confidence the proportion of people who own a desktop computer. A previous study showed that 40% of those interviewed owned a desktop. The researcher wishes to be accurate within 2% of the true proportion.

From the standard normal table,

95% confidence interval

Confidence interval is 95%

95% = 95/100 = 0.95

\alpha = 1 - Confidence interval = 1-0.95 = 0.05

\alpha/2 = 0.05 / 2

= 0.025

Z\alpha/2 = Z0.025 = 1.96

the z-critical value at 95% confidence is 1.96

The information represent

E = 2% = 2/100 = 0.02

\hat{p} = 40% = 40/100 = 0.40

Find the minimum sample size

sample size = n = (Z\alpha/2 / E)2 * \hat{p} (1-\hat{p})

n = (1.96/ 0.02)2 * 0.40(1-0.40)

n = 9604 *0.24

n = 2304.96

The minimum sample size is 2305

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.